The seventh term of an A.P is 15 and the fourth term is 9. Find the first term and common difference
1
Expert's answer
2020-04-08T12:12:31-0400

According to the formula of A.P.:

a7 = a1 + (7 - 1)d and a4 = a1 + (4 - 1)d,

15 = a1 + 6d and 9 = a1 + 3d,

a1 = 15 - 6d = 9 - 3d,

3d = 6,

d = 2 is the common difference.

a1 = 9 - 3×2 = 3 is the first term.